# Large Language Model-driven Meta-structure Discovery in Heterogeneous Information Network

### v1

@@ 5.2. Deep Learning on HIN
The success of graph neural networks introduces revolutionary deep learning techniques into HIN modeling (Wu et al., 2020). Metapath2Vec (Dong et al., 2017) proposed to learn deep representations for nodes in HIN via meta-path-guided random walks. Attention mechanism was later introduced to learn more expressive representations for HIN, which were proven effective for link prediction (Xu et al., 2019) and node classification (Wang et al., 2019). Extensive follow-up research was devoted to designing more effective heterogeneous GNN frameworks (Li et al., 2021a; Fu et al., 2020). Besides, considerable research efforts were drawn to replace handcrafted meta-structures with automatic search. GEMS (Han et al., 2020) proposed to combine heterogeneous GNN with evolutionary algorithms to identify meta-structures and learn deep neural networks simultaneously. Several deep reinforcement learning models and neural architecture search models are also proposed to jointly optimize the meta-structures with heterogeneous GNN (Ding et al., 2021; Peng et al., 2021; Ning et al., 2022; Li et al., 2023; Wan et al., 2020; Li et al., 2021b). However, previous automatic meta-structure design method solely focused on prediction performance, often yielding complex and difficult to explain meta-structures. To the best of our knowledge, our study is the first to harness the semantic reasoning capability of LLMs for automatic meta-structure design. Our model can discover meta-structures that not only show high prediction performance, but also can be adequately explained with natural languages.
@@ 5.3. LLM for Graph Learning
The recent advancement of LLMs shows general capabilities that go beyond natural language tasks. Researchers on graph learning are particularly interested in their ability to graph reasoning tasks (Guo et al., 2023a). For example, there are several recent attempts to use LLMs to enhance node features or employ them as standalone graph predictors (Chen et al., 2023). The research communities are also actively discussing the perspective for developing large graph models (Zhang et al., 2023; Liu et al., 2023). Recent research found LLMs exhibited preliminary ability for graph learning tasks, such as detecting connectivity and cycles, performing topological sort, and emulating GNNs (Wang et al., 2023). Besides, LLMs can also effectively perform reasoning tasks on knowledge graph (Sun et al., 2023). To better align graph problems with LLM, several recent works propose various methods to encode the geometric structure and node features of graph problems (Zhao et al., 2023; Fatemi et al., 2023). On top of these textual encodings, researchers have explored the possibility of replacing GNNs with LLM reasoning (Ye et al., 2023) and performing instruction tuning for graph tasks (Tang et al., 2023). In this paper, we fundamentally extend LLM reasoning to HIN meta-structure discovery. Specifically, we propose a novel meta-structure encoding method, which effectively boosts LLM’s reasoning capability on HINs.
@@ 5.4. LLM for Pattern Discoveries
The scaled-up language models have emerged reasoning capability for general tasks (Wei et al., 2022b), including commonsense reasoning, logical reasoning, and mathematics reasoning. Optimized prompting routines like chain-of-thought (CoT) (Wei et al., 2022b) and tree-of-thoughts (ToT) (Yao et al., 2023) can effectively improve the scaling curve of LLM’s reasoning capability. As a result, recent research showed that LLM can be leveraged to identify novel patterns and feasible solutions in a large space. For example, FunSearch was proposed for mathematical algorithms discovery (Romera-Paredes et al., 2023). Besides, previous works also designed novel LLM-driven algorithms for evolutionary search (Guo et al., 2023b), reinforcement learning (Shinn et al., 2023), and hyper-parameter optimization (Liu et al., 2024). In this paper, we propose a novel framework to harness the reasoning power of LLM for meta-structure discovery. Our framework equips LLM with enhanced capability to understand the semantic meaning of meta-structures and search for promising candidates.
@@ 6. Conclusion
In this work, we propose a novel framework, ReStruct, that fuses the power of LLMs with evolutionary algorithms to facilitate automatic meta-structure discovery across diverse HINs. Extensive experiments demonstrate that our model excels in uncovering previously undiscovered meta-structures, thereby significantly enhancing downstream model performance compared to state-of-the-art baselines. Notably, a pilot study involving human participants confirms that our model provides a substantial level of interpretability regarding the identified meta-structures. For future work, we will explore the feasibility of finetuning local models to mitigate network communication costs associated with API calls.
@@ Appendix A Statistics of Datasets
Figure 9. Statistics of datasets for node classification
Figure 10. Statistics of datasets for recommendation
Appendix B Questionnaire for Meta-Structure Explainability
Please find the questionnaire in the next few pages.

### v2

@@ 5.2. Deep Learning on HINs
The success of graph neural networks introduces revolutionary deep learning techniques into HIN modeling (Wu et al., 2020). Metapath2Vec (Dong et al., 2017) proposed to learn deep representations for nodes via meta-path-guided random walks. Attention mechanism was later introduced to learn more expressive representations for HINs, proving effective for link prediction (Xu et al., 2019) and node classification (Wang et al., 2019). Subsequent research endeavors focused on designing more effective heterogeneous GNN frameworks (Li et al., 2021a; Fu et al., 2020). Besides, considerable research efforts were drawn to replace handcrafted meta-structures with automatic search. GEMS (Han et al., 2020) proposed to combine heterogeneous GNN with evolutionary algorithms to identify meta-structures and learn deep neural networks simultaneously. Several deep reinforcement learning models and neural architecture search models are also proposed to jointly optimize the meta-structures with heterogeneous GNN (Ding et al., 2021; Peng et al., 2021; Ning et al., 2022; Li et al., 2023; Wan et al., 2020; Li et al., 2021b). However, previous automatic meta-structure design method solely focused on prediction performance, often yielding complex and difficult to explain meta-structures. To the best of our knowledge, our study is the first to harness the semantic reasoning capability of LLMs for automatic meta-structure design. Our model can discover meta-structures that not only show high prediction performance, but also can be adequately explained with natural languages.
@@ 5.3. LLM for Graph Learning
LLMs have demonstrated general capabilities beyond natural language tasks, attracting graph learning researchers who are particularly interested in leveraging their ability for graph reasoning tasks (Guo et al., 2023a). Several attempts have been made to enhance node features using LLMs or employ them as standalone graph predictors (Chen et al., 2023). The research community is also actively discussing the perspective of developing large graph models (Zhang et al., 2023; Liu et al., 2023). Recent research has found that LLMs exhibit certain ability for graph tasks such as detecting connectivity and cycles, performing topological sort, and emulating GNNs (Wang et al., 2023). Besides, LLMs can effectively perform reasoning on knowledge graphs (Sun et al., 2023b). To better align graph problems with LLMs, recent works propose various methods to encode the geometric structure and node features of graph problems (Zhao et al., 2023; Fatemi et al., 2023). With these encodings, researchers have explored the possibility of replacing GNNs with LLM reasoning (Ye et al., 2023) and performing instruction tuning for graph tasks (Tang et al., 2023). In this paper, we fundamentally extend LLM reasoning to HIN meta-structure discovery. Specifically, we propose a novel meta-structure encoding method, which effectively boosts LLMs’ reasoning capability on HINs.
@@ 5.4. LLM for Pattern Discoveries
The scaled-up language models have emerged reasoning capability for general tasks (Wei et al., 2022b), including commonsense reasoning, logical reasoning, and mathematics reasoning. With the help of optimized prompting routines such as chain-of-thought (CoT) (Wei et al., 2022b) prompting and tree-of-thoughts (ToT) (Yao et al., 2023) prompting, the scaling curve of LLMs’ reasoning capability can be further effectively improved. As a result, recent research has shown that LLMs can be leveraged to identify novel patterns and feasible solutions in large problem spaces. For example, FunSearch was proposed to discover algorithm programs for solving mathematical problems (Romera-Paredes et al., 2023). Previous works also designed LLM-driven algorithms for evolutionary search (Guo et al., 2023b), reinforcement learning (Shinn et al., 2023), and hyper-parameter optimization (Liu et al., 2024). In this paper, we propose a novel framework to harness the reasoning power of LLM for meta-structure discovery. Our framework equips LLMs with enhanced capability to understand the semantic meaning of meta-structures and search for promising candidates.
@@ 6. Conclusion
In this work, we propose a novel framework, ReStruct, that fuses the power of LLMs with evolutionary algorithms to facilitate automatic meta-structure discovery across diverse HINs. On both recommendation and node classification tasks, extensive experiments demonstrate that ReStruct excels in uncovering previously undiscovered meta-structures, thereby significantly enhancing downstream model performance compared to a set of state-of-the-art baselines. Notably, a user study involving human participants confirms that ReStruct substantially outperforms baseline methods in terms of the comprehensibility of discovered meta-structures and usefulness of generated explanations. For future work, we will explore the feasibility of finetuning local models to mitigate network communication costs associated with API calls.
7. Acknowledgement
This work was supported in part by the National Key Research and Development Program of China under Grant 23IAA02114, the Guangzhou Municipal Nansha District Science and Technology Bureau under Contract No.2022ZD012, and the National Natural Science Foundation of China under Grant U23B2030 and Grant U22B2057.
@@ Appendix A Statistics of Datasets
Table 5. Statistics of datasets for node classification.
Dataset ACM IMDB DBLP OAG-NN
Node
types
Author (A)
Paper (P)
Subject (S)
Movie (M)
Actor (A)
Director (D)
Author (A)
Paper (P)
Conference (C)
Paper (P)
Author (A)
Affiliation (I)
Field (F)
Edge
types
A-P, P-A,
P-S, S-P
M-D, D-M,
M-A, A-M
A-P, P-A,
P-C, C-P
P-P,
P-A, A-P,
P-F, F-P,
A-I, I-A
# Nodes 8,994 12,624 18,405 64,203
# Edges 25,922 37,288 67,946 403,974
# Classes 3 3 4 8
# Training 600 300 800 2,334
# Validation 300 300 400 778
# Testing 2,125 2,339 2,857 778
Table 6. Statistics of datasets for recommendation.
Dataset Relations (S-T) # S # T # S-T
Yelp User-Business (U-B) 16,239 14,284 84,993
User-User (U-U) 16,239 16,239 158,590
User-Compliment (U-O) 16,239 11 76,875
Business-Category (B-A) 14,284 511 40,009
Business-City (B-I) 14,284 47 14,267
Douban User-Movie (U-M) 13,367 12,677 500,515
User-User (U-U) 13,367 13,367 4,085
User-Group (U-G) 13,367 2,753 570,047
Movie-Actor (M-A) 12,677 6,311 33,587
Movie-Director (M-D) 12,677 2,449 11,276
Movie-Type (M-T) 1,2677 38 27,668
Amazon User-Item (U-I) 6,170 2,753 86,191
Item-View (I-V) 2,753 3,857 5,694
Item-Category (I-C) 2,753 22 5,508
Item-Brand (I-B) 2,753 334 2,753
LastFM User-Artist (U-A) 1,892 17,632 46,417
User-User (U-U) 1,892 1,892 25,434
Artist-Tag (A-T) 17,632 9,718 108,437
Appendix B Emergence of Occam’s Razor Phenomenon
In our user study, we demonstrate that ReStruct finds meta-structures that are more comprehensible to human researchers (Figure 6), implying a tendency to avoid overcomplicated meta-structures. To provide further evidence, we add an experiment on the Yelp recommendation task, asking the LLM to select between meta-structure pairs with near-equal performance on the validation set but varying structural complexity (#nodes and #edges), which are visualized in Figure 8. As shown in Table 7, LLM consistently prefers meta-structures with lower structural complexity, citing reasons such as ”simplicity,” ”generalizability,” and even invoking ”Occam’s razor”. Therefore, LLMs possess human-like preferences for simpler and more comprehensible meta-structures that mitigates overfitting.
Figure 8. Visualization of meta-structure pairs for comparison.
[…]
Visualization of meta-structure pairs for comparison.
Table 7. LLM selection between meta-structure pairs with near-equal validation performances: three examples.
Pair Chosen Index Reasons
1 (Fig. 8A) 1 “- Complexity vs. Performance: The second meta-structure has a more complex path, suggesting potentially richer semantics being captured. However, if both meta-structures show almost equal performance, the additional complexity might not translate into a significant performance gain. In such cases, it is generally better to opt for the simpler model (Occam’s razor), which is why Meta-structure 1 is preferred. - Overfitting Risks: The more complex second meta-structure might be more prone to overfitting, as it captures a more specific set of interactions. The simpler Meta-structure 1 might generalize better to unseen data. - Interpretability: Meta-structure 1 is easier to interpret because it has a shorter path and fewer node interactions. This can be beneficial for understanding why certain recommendations are being made, which is important for transparency and trust in the recommendation system. - Computational Efficiency: With shorter paths and fewer relations, Meta-structure 1 is likely to be more computationally efficient for both training and inference, which can be important in practical, large-scale applications. - Meta-path Relevance: The paths in Meta-structure 1 adequately capture the relevant user-business interactions via a business’s location. Adding more nodes and relations in the second meta-structure might not contribute much to capturing the essential features for the recommendation task. In conclusion, the first meta-structure is selected as it achieves a good balance between simplicity, performance, and interpretable results while potentially being more computationally efficient and generalizable.”
2 (Fig. 8B) 1 “- Path Simplicity: Meta-structure 1 offers a more straightforward path from ’U’ (User) to ’B’ (Local Business) through ’A’ (Business Category) and ’I’ (Business City), with minimal backtracking or redundant connections. The path follows a logical progression from user preference to business through the context of category and location. - Computational Efficiency: The simpler meta-structure 1 should, in theory, be computationally less intensive because it involves fewer steps and connections. This could be advantageous when scaling the recommendation system and processing a large dataset. - Interpretability: Meta-structure 1 is easier to interpret. In the context of a recommendation system, a clear logical flow from user to business via category and city is likely to align with a user’s thought process when seeking recommendations, enhancing the user experience. - Relevance of Categories and Location: By going through ’A’ (Business Category) and ’I’ (Business City), Meta-structure 1 reflects important aspects of businesses that users consider while making choices—what kind of service or product a business offers (category) and where it is located (city). While Meta-structure 2 includes a potentially rich and complex connection pattern that might capture more nuanced relationships, the balance between complexity and utility must be considered, particularly if both structures have shown similar performance on the validation set. Given these considerations, the simpler, more interpretable, and likely more efficient Meta-structure 1 would likely be the optimal choice for the recommendation model.”
3 (Fig. 8C) 1 “- **Simplicity and Interpretability**: The first meta-structure is less complex as it contains a shorter sequence of nodes, potentially making it easier to interpret and explain how the recommendation is generated. A simpler model is often preferred when performances are almost equal because it can be more transparent, and it can generalize better, avoiding overfitting the data. - **Computational Efficiency**: Fewer node hops suggest lower computational costs in terms of executing the meta-path. This can be important for scaling the recommendation model to a large number of users or in an environment where computational resources are constrained. - **Redundancy Avoidance**: The second meta-structure involves a repeated sequence of ’Business Category’ to ’Local Business’ (’A’ to ’B’) hops, which may introduce redundancy that does not necessarily contribute to predictive performance. Such over-specified meta-structures can sometimes lead to overfitting, where the performance on the validation set is good, but the generalization to unseen data may suffer. - **Logical Coherence**: The first meta-structure provides a more direct flow of information from ’User’ to ’Local Business’ through their categories and mutual connections to other users. This logical flow is more coherent compared to the second meta-structure, which seems to loop back through ’Business Category’ without a clear additional benefit. - **Real-World Significance**: From a real-world perspective, the first meta-structure captures a user’s preference through the categories of businesses they have visited and the influence of their friends’ preferences for the same business categories. This could be a robust basis for recommendations without the need for the extra category-business-category loop in the second meta-structure. Given these reasons, meta-structure 1 seems to be the optimal choice for improving the recommendation model on this Yelp HIN.”
Appendix C Evidence of opportunity-risk tradeoff during candidate selection
To check whether there is an overreliance on confidence values during candidate selection, we add an experiment to compare the confidence values between chosen and unchosen meta-structure candidates. As shown in Figure 9, ReStruct does not adhere strictly to the candidates with the highest confidence values. Take Experiment #14 as an example: although Candidate 00 and $5$ have higher confidence values, ReStruct decides to choose Candidate $3$, likely due to its higher predicted performance via effective semantic analysis. This highlights ReStruct’s attempts to balance between opportunities and risks, when predicted performances and confidence values become trade-offs.
Figure 9. Comparison of confidence values between chosen and unchosen candidates.
[…]
Comparison of confidence values between chosen and unchosen candidates.